OpenStudy (anonymous):

If f(x) varies directly with x, and f(x) = 10 when x = 4, write the direct linear variation equation.

OpenStudy (anonymous):

answers will come in a sec

OpenStudy (anonymous):

a. \[f(x)=10x\] b. \[f(x)=4x\] c. (5/2 is a fraction) \[f(x)=5/2x\] d. (2/5 is a fraction) \[f(x)=2/5x\]